OpenStudy (anonymous):

find the values of the six trigonometric functions of an angle Q in standard position whose terminal side passes through (1, -2)

OpenStudy (mertsj):

If x = 1 and y = -2, the angle is in the 4th quadrant and the hypotenuse is sqrt5 So sin = -2sqrt5/5, cosine = sqrt5/5, tangent = -2, cotangent = -1/2 sec = sqrt5 csc = -sqrt5/2

OpenStudy (mertsj):